OpenStudy (anonymous):

Help please will give a medal. - 1 over (w-3) = - 3 over (2w - 6) = 3 solve for w ** I came up with no solution but that is wrong **

OpenStudy (anonymous):

unreadable i am afraid

OpenStudy (anonymous):

same to me

OpenStudy (anonymous):

@satellite73 I was still typing it when you looked at it can you please look at it now.

OpenStudy (anonymous):

can you help me : The distance between each of the parts of a personal letter that has been handwritten is inch.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

let me see if i can read it

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

\[\frac{-1}{w-3}=\frac{-3}{2w-6}\] is that it?

OpenStudy (anonymous):

yes but there is also a -3 after those two parts

OpenStudy (anonymous):

can you post a screen shot?

OpenStudy (anonymous):

\[\frac{-1}{w-3}=\frac{-3}{2w-6}-3\]

OpenStudy (anonymous):

yes that is it

OpenStudy (anonymous):

ok lets add \(\frac{3}{2w-6}\) to both sides and get \[\frac{3}{2w-6}-\frac{1}{w-3}=-3\]

OpenStudy (anonymous):

on the left the common denominator is \(2w-6\) so we and do this \[\frac{3}{2w-6}-\frac{2}{2w-6}=\frac{1}{2w-6}\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

now you can solve \[\frac{1}{2w-6}=-3\] by writing \[1=-3(2w-6)\] and solve for \(w\)

OpenStudy (anonymous):

w - 17 over 6 (17/6)

OpenStudy (anonymous):

thank you so much that makes sense now.

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

\[1=-3(2w-6)\\ 1=-6x+18\\ -17=-6x\\ x=\frac{17}{6}\] just checking

OpenStudy (anonymous):

im still confused -_-